OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line with slope and containing the midpoint of the segment whose endpoints are (2, -3) and (-6, 5).

OpenStudy (anonymous):

The equation of a line is: \[y = mx + b\] where \[m\] is the slope and \[b\] is the y-intercept. It looks like you're missing part of the question where they specify the slope of the line you are looking for, but once you have it you would solve the question as follows: 1) Find the midpoint of the line segment, using\[(x1 + x2)/2\] for the X-coordinate, and \[(y1 + y2)/2\] for the y-coordinate. 2) Using our formula for a line, fill in the \[y, x, & m\] variables & solve for b, or alternatively, use the formula \[y-y0 = m(x-x0)\] where x0 & y1 are the midpoint mentioned above, and m is the slope.

OpenStudy (anonymous):

thx man helped alot :)

OpenStudy (anonymous):

No worries!